OpenStudy (anonymous):

14x - (x2 + 8x) = btw, my names Karla :D*

OpenStudy (anonymous):

14x - (x^2 + 8x) = 0 14x - x^2 -8x = 0 6x - x^2 = 0 x^2 - 6x = 0 x (x-6) = 0 \[x _{1} = 0\]\[x _{2} = 6\] You are welcome.

OpenStudy (anonymous):

thank you :)